# Beyond MMD: Evaluating Graph Generative Models with Geometric Deep Learning

### v1

@@ IV Experiments
This section presents the experimental setup and results of the classification task within our RGM framework, along with the training, generation, and evaluation procedures for the two selected GGMs: GRAN and EDGE. The motivation behind selecting these two models lies in their distinct capabilities and generative approaches, which complement each other in capturing different aspects of graph structures. GRAN is an auto-regressive model that has demonstrated superior performance compared to GraphRNN [you2018graphrnngeneratingrealisticgraphs] in generating realistic graphs, particularly for datasets with complex structural dependencies. On the other hand, EDGE is a diffusion-based model that leverages iterative refinement to generate graphs, making it well-suited for capturing fine-grained details and global graph properties.
@@ IV-A Experimental Setup
The experiments were conducted on a custom dataset drawn from the publicly available graph collection at the Network Data Repository. The dataset consists of graphs from five distinct classes: Biological, Connectome, Infrastructure, Internet, and Social. To further assess the robustness of the model and extend its applicability across a broader range of graph types, synthetic graphs were incorporated into the dataset. These synthetic graphs were generated using five well-known graph generation models: the Barabási-Albert model, the Erdős-Rényi model, the Stochastic Block Model, the Lancichinetti-Fortunato-Radicchi model, and the nonuniform Popularity-Similarity-Optimization model. A total of $300$ graphs were generated for each model, featuring diverse statistics, using NetworkX [hagberg2008exploring] and iGraph [Csardi2006-qq] libraries. This expanded dataset, encompassing $10$ classes in total, provides a comprehensive foundation for evaluating the performance of graph generative models across a wide range of applications and domains.
@@ IV-B Representation-aware Graph-generation Model evaluation (RGM)
The input node features are computed using a set of topological features that are extracted from the graphs in the dataset. Specifically, we use four input topological features that capture multiscale information about the graph structure: the node degree, a chi-square statistic computed over the node’s neighborhood, the local clustering coefficient, and the k-core number. While the node degree and chi-square statistic capture local connectivity patterns, the clustering coefficient and k-core number provide information about tightly connected groups of nodes. For each graph, the node features are concatenated into a single feature vector, which is then fed into the GNN layers that propagate and aggregate them iteratively during the message-passing process. For instance, when propagated, the node degree can be used to infer assortativity and heterogeneity in the graph structure, while the chi-square statistic can help identify local patterns and outliers. The local clustering coefficient captures the tendency of nodes to cluster together locally, indicating the presence of tightly-knit communities or groups within the graph, while the k-core number provides information about the core structure of the graph. The features, computed using graph-tool [peixoto_graph-tool_2014], are leveraged to enrich the node representations and improve the model’s ability to capture the underlying structure of the graphs, as shown in [grassia2021machine].
The model was trained using a supervised approach, where the input consists of triplets of graphs: an anchor graph, a positive graph (same class as the anchor), and a negative graph (different class). The framework is represented in Figure 1. The training process was performed over $64\%$ of the dataset using a Triplet Margin Loss function, which encourages the model to learn embeddings that are closer for similar graphs and farther apart for dissimilar graphs, using the AdamW optimizer with a learning rate of $0.003$ and a weight decay of $5\times 10^{-3}$; early stopping was used based on the validation loss. The validation set consisted of $16\%$ of the total dataset, while the remaining $20\%$ of the graphs were reserved for testing. Test graphs are classified into their respective classes using a balanced anchor selection strategy based on a dynamic $k$-NN approach. Specifically, each generated graph is assigned to a class via a $k$-nearest neighbor classifier, which compares its embedding to those of the training anchors and selects the class of the closest ones. The number of neighbors $k_{i}$ used for each class $i$ is determined dynamically as $k_{i}=\sqrt{|\mathrm{anchors}_{i}|}$, where $|\mathrm{anchors}_{i}|$ is the number of training anchors for that class, following the recommendation in [DBLP:journals/corr/HassanatAAA14].
@@ IV-C Graph Generation: Assessment with RGM
Here, we computed the MMD values for the following metrics: degree, clustering coefficient, Orbits, spectral, and Neighborhood Subgraph Pairwise Distance Kernel (NSPDK). In particular, the degree and the clustering coefficient distributions are local properties of the nodes that capture the number of connections and the tendency of the node’s neighbors to connect with each other, respectively. The Orbits metric, computed using the Orca algorithm [hovcevar2014combinatorial], captures the number of node- and link-orbits (i.e., the number of distinct ways a node can be connected to its neighbors) of 4- and 5-node graphlets in the network, providing a measure of the local structure around each node. The spectral metric captures the eigenvalues of the adjacency matrix, which can provide insights into the global structure of the graph. Finally, the NSPDK (Neighborhood Subgraph Pairwise Distance Kernel) [costa2010fast] metric captures the distance between pairs of nodes in the graph based on their neighborhood structure, providing a measure of the similarity between different parts of the graph.
@@
In the MMD evaluation Tables V and VI, the computed MMD values between the reference and generated datasets show mixed results even after extensive training, with most metrics indicating good alignment while others reveal significant discrepancies [you2018graphrnngeneratingrealisticgraphs, liao2020efficientgraphgenerationgraph]. For instance, clustering coefficient MMD values are particularly high for Internet graphs (0.780 for GRAN, 0.720 for EDGE) and Connectome graphs (0.874 for EDGE), while spectral properties show substantial deviations, especially for Connectome graphs (1.146 for EDGE, 0.747 for GRAN). In any case, despite most low MMD values (e.g., <0.05<0.05 for several NSPDK and Orbits metrics), our RGM framework struggles to accurately classify the generated graphs. Based on the previously reported results, we can assume that our model captures relevant structural differences. Its poor performance in assigning the correct domain to generated graphs suggests that they still lack important characteristics of the original data, highlighting serious limitations of current state-of-the-art GGMs. This discrepancy also highlights a key limitation of current usage of MMD: low values across certain statistics do not necessarily imply high structural fidelity or downstream utility. Although individual metrics may appear well aligned, the overall similarity remains inadequate, as also shown by the statistical comparison of a wide set of network science metrics. In fact, by looking at Figures A.1, A.2 and A.3 in the Appendix, we can notice that there are substantial deviations in key topological properties when compared to the original graphs (i.e., the baseline used to train the GGMs). For instance, both EDGE and GRAN struggle to accurately replicate assortativity, average clustering coefficient and transitivity, affecting the homophily and the formation of tightly-knit groups in the network. The density, diameter and average path length are also misaligned, especially for GRAN. The k-core analysis exhibits a significant mismatch in the Internet class for GRAN, and both GRAN and EDGE fail to capture the expected k-core structure in the Biological and Social classes.
@@ V Conclusion
In this work, we present a novel framework for evaluating the generative capabilities of Graph Generative Models (GGMs) using Geometric Deep Learning (GDL) techniques. Our framework, RGM (Representation-aware Graph-generation Model evaluation), is based on a Siamese Graph Neural Network trained to recognize similar structural patterns in graphs and is capable of classifying graphs into their respective domains based on their topological structure with high accuracy.
@@ VI Acknowledgment
VII Biography Section
[Uncaptioned image] Salvatore Romano is a PhD student enrolled in the National PhD in Artificial Intelligence, XXXIX cycle, organized by University Campus Bio-Medico of Rome (Rome, Italy) and affiliated with University of Catania (Catania, Italy). He received his MSc in Data Science from the University of Catania. His research interests span from geometric deep learning to graph generation.
[Uncaptioned image] Marco Grassia is an Assistant Professor at the Department of Electrical, Electronic and Computer Engineering (DIEEI) of the University of Catania (Catania, Italy). He received his PhD in Computer Science from the University of Catania in 2018. His research interests include network science, geometric deep learning and their applications to complex systems.
[Uncaptioned image] Giuseppe Mangioni is associate professor in Department of Electrical Electronics and Computer Engineering at the University of Catania. He received the degree in Computer Engineering (1995) and the Ph.D. degree (2000) at the University of Catania. Currently he teaches algorithms, internet and security and OOP at the University of Catania. His research interests include network science, machine learning and geometric deep learning. He is co-initiator and chair of the steering committee of the Int. Conference on Complex Networks (CompleNet).

### v2

@@ IV Experiments
This section presents the experimental setup and results of the classification task within our RGM framework, along with the training, generation, and evaluation procedures for the two selected GGMs: GRAN and EDGE. The motivation behind selecting these two models lies in their distinct capabilities and generative approaches, which complement each other in capturing different aspects of graph structures. GRAN is an auto-regressive model that has demonstrated superior performance compared to GraphRNN [43] in generating realistic graphs, particularly for datasets with complex structural dependencies. On the other hand, EDGE is a diffusion-based model that leverages iterative refinement to generate graphs, making it well-suited for capturing fine-grained details and global graph properties.
@@ IV-A Experimental Setup
The experiments were conducted on a custom dataset drawn from the publicly available graph collection at the Network Data Repository. The dataset consists of graphs from five distinct classes: Biological, Connectome, Infrastructure, Internet, and Social. To further assess the robustness of the model and extend its applicability across a broader range of graph types, synthetic graphs were incorporated into the dataset. These synthetic graphs were generated using five well-known graph generation models: the Barabási-Albert model, the Erdős-Rényi model, the Stochastic Block Model, the Lancichinetti-Fortunato-Radicchi model, and the nonuniform Popularity-Similarity-Optimization model. A total of $300$ graphs were generated for each model, featuring diverse statistics, using NetworkX [18] and iGraph [10] libraries. This expanded dataset, encompassing $10$ classes in total, provides a comprehensive foundation for evaluating the performance of graph generative models across a wide range of applications and domains.
@@ IV-B Representation-aware Graph-generation Model evaluation (RGM)
The input node features are computed using a set of topological features that are extracted from the graphs in the dataset. Specifically, we use four input topological features that capture multiscale information about the graph structure: the node degree, a chi-square statistic computed over the node’s neighborhood, the local clustering coefficient, and the k-core number. While the node degree and chi-square statistic capture local connectivity patterns, the clustering coefficient and k-core number provide information about tightly connected groups of nodes. For each graph, the node features are concatenated into a single feature vector, which is then fed into the GNN layers that propagate and aggregate them iteratively during the message-passing process. For instance, when propagated, the node degree can be used to infer assortativity and heterogeneity in the graph structure, while the chi-square statistic can help identify local patterns and outliers. The local clustering coefficient captures the tendency of nodes to cluster together locally, indicating the presence of tightly-knit communities or groups within the graph, while the k-core number provides information about the core structure of the graph. The features, computed using graph-tool [37], are leveraged to enrich the node representations and improve the model’s ability to capture the underlying structure of the graphs, as shown in [13].
The model was trained using a supervised approach, where the input consists of triplets of graphs: an anchor graph, a positive graph (same class as the anchor), and a negative graph (different class). The framework is represented in Figure 1. The training process was performed over $64\%$ of the dataset using a Triplet Margin Loss function, which encourages the model to learn embeddings that are closer for similar graphs and farther apart for dissimilar graphs, using the AdamW optimizer with a learning rate of $0.003$ and a weight decay of $5\times 10^{-3}$; early stopping was used based on the validation loss. The validation set consisted of $16\%$ of the total dataset, while the remaining $20\%$ of the graphs were reserved for testing. Test graphs are classified into their respective classes using a balanced anchor selection strategy based on a dynamic $k$-NN approach. Specifically, each generated graph is assigned to a class via a $k$-nearest neighbor classifier, which compares its embedding to those of the training anchors and selects the class of the closest ones. The number of neighbors $k_{i}$ used for each class $i$ is determined dynamically as $k_{i}=\sqrt{|\mathrm{anchors}_{i}|}$, where $|\mathrm{anchors}_{i}|$ is the number of training anchors for that class, following the recommendation in [21].
@@ IV-C Graph Generation: Assessment with RGM
Here, we computed the MMD values for the following metrics: degree, clustering coefficient, Orbits, spectral, and Neighborhood Subgraph Pairwise Distance Kernel (NSPDK). In particular, the degree and the clustering coefficient distributions are local properties of the nodes that capture the number of connections and the tendency of the node’s neighbors to connect with each other, respectively. The Orbits metric, computed using the Orca algorithm [22], captures the number of node- and link-orbits (i.e., the number of distinct ways a node can be connected to its neighbors) of 4- and 5-node graphlets in the network, providing a measure of the local structure around each node. The spectral metric captures the eigenvalues of the adjacency matrix, which can provide insights into the global structure of the graph. Finally, the NSPDK (Neighborhood Subgraph Pairwise Distance Kernel) [9] metric captures the distance between pairs of nodes in the graph based on their neighborhood structure, providing a measure of the similarity between different parts of the graph.
@@
In the MMD evaluation Tables V and VI, the computed MMD values between the reference and generated datasets show mixed results even after extensive training, with most metrics indicating good alignment while others reveal significant discrepancies [43, 30]. For instance, clustering coefficient MMD values are particularly high for Internet graphs (0.780 for GRAN, 0.720 for EDGE) and Connectome graphs (0.874 for EDGE), while spectral properties show substantial deviations, especially for Connectome graphs (1.146 for EDGE, 0.747 for GRAN). In any case, despite most low MMD values (e.g., <0.05<0.05 for several NSPDK and Orbits metrics), our RGM framework is unable to accurately classify the generated graphs. Based on the previously reported results, we can assume that our model captures relevant structural differences. Its poor performance in assigning the correct domain to generated graphs suggests that they still lack important characteristics of the original data, highlighting serious limitations of current state-of-the-art GGMs. This discrepancy also highlights a key limitation of current usage of MMD: low values across certain statistics do not necessarily imply high structural fidelity or downstream utility. Although individual metrics may appear well aligned, the overall similarity remains inadequate, as also shown by the statistical comparison of a wide set of network science metrics. In fact, by looking at Figures A.1, A.2 and A.3 in the Appendix, we can notice that there are substantial deviations in key topological properties when compared to the original graphs (i.e., the baseline used to train the GGMs). For instance, both EDGE and GRAN struggle to accurately replicate assortativity, average clustering coefficient and transitivity, affecting the homophily and the formation of tightly-knit groups in the network. The density, diameter and average path length are also misaligned, especially for GRAN. The k-core analysis exhibits a significant mismatch in the Internet class for GRAN, and both GRAN and EDGE fail to capture the expected k-core structure in the Biological and Social classes.
@@ V Conclusion
In this work, we present a novel framework for evaluating the generative capabilities of Graph Generative Models (GGMs) using Geometric Deep Learning (GDL) techniques. Our framework, RGM (Representation-aware Graph-generation Model evaluation), is based on a Siamese Graph Neural Network trained to recognize similar structural patterns in graphs, and is capable of classifying graphs into their respective domains based on their topological structure with high accuracy.